OpenStudy (anonymous):

Find the first four terms of the sequence represented by the expression 8(1/2)n

OpenStudy (anonymous):

\(8\times \left(\frac{1}{2}\right)^n\)

OpenStudy (anonymous):

replace \(n\) by 1 get \(8\times \left(\frac{1}{2}\right)^1=8\times \frac{1}{2}=4\)

OpenStudy (anonymous):

now for the next one, you can replace \(n\) by 2, or more simply take half of 4 third one is half of 2 etc

OpenStudy (anonymous):

ok thanks!! @satellite73

OpenStudy (anonymous):

yw what a random collection of topics you have what course is this?

OpenStudy (anonymous):

@satellite73 its a bunch mixed together ik it is really weird